OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation x^2/81+y^2/225=1 @agent0smith

OpenStudy (agent0smith):

Center is (0, 0) foci will be (0, c) and (0, -c) where c^2 = b^2 - a^2 and b^2 = 225 and a^2 = 81

OpenStudy (anonymous):

Center: (0, 0); Vertices: (0, -15), (0, 15); Foci: (0, -12), (0, 12)

OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 2x2 + 8y2 = 16.

OpenStudy (agent0smith):

Divide everything by 16, then it's almost the same as above, but this one is horizontal, so foci (c, 0) and (-c, 0) where c^2 = a^2 - b^2

OpenStudy (anonymous):

(0,-sqrt 6) (0,sqr 6)

OpenStudy (agent0smith):

close but see above post

OpenStudy (anonymous):

oh wait -sqt6,0 and sqrt6,0

OpenStudy (anonymous):

no/?

OpenStudy (agent0smith):

Yes